You need to prepare an acetate buffer of pH 5.23 from a 0.815 M acetic acid solution and a 2.26 M KOH solution. If you have 680 mL of the acetic acid solution, how many milliliters of the KOH solution do you need to add to make a buffer of pH 5.23? The p?apKa of acetic acid is 4.76.
Answer #1

millimoles of acetic acid = 680 x 0.815 = 554.2

millimoles of KOH = 2.26 V

CH3COOH   +    KOH    -------------> CH3COOK   + H2O

554.2                 2.26 V                          0                    0

554.2-2.26V         0                              2.26 V

pH = pKa + log [salt/ acid]

5.23 = 4.76 + log [2.26 V /554.2-2.26V ]

[2.26 V /554.2-2.26V ] = 2.951

2.26 V = 1635.56 - 6.67 V

V = 183.2 mL

volume of KOH = 183 mL